Which of the following can be the sides of a right triangle?
 2 cm, 2 cm, 5 cm.
In the case of right-angled triangles, identify the right angles

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Let a = 2 cm, b = 2 cm, c = 5 cm
Let us assume the largest value is the hypotenuse side i.e. c = 5 cm. Then, by Pythagoras theorem,
c^2=a^2+b^2
5^2=2^2+2^2
25 = 4 + 4
25≠8
The sum of square of two side of triangle is not equal to the square of third side,
∴The given triangle is not right-angled triangle.

Thanks